OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (−4, 8) and (−3, −6)

OpenStudy (anonymous):

The slope of the line, or gradient, or \(m\), passing throught two points \((x_1,y_1) and (x_2,y_2)\)can be found with this formula: \[m=\frac{y_2-y_1}{x_2-x_1}\]

hero (hero):

Insert each point into y = mx + b as follows: 8 = 4m + b -6 = -3m + b Isolate b in each equation as follows: b = 8 - 4m b = 3m - 6 Set b = b and solve for m, the slope: 8 - 4m = 3m - 6 8 + 6 = 3m + 4m 14 = 7m 14/7 = m 2 = m Therefore, the slope of the line is 2